OpenStudy (anonymous):

integration by parts question. what is the indefinite integral of sqrt(t)ln(t)dt? I got (2/3)(t^(3/2))(ln(t) - 1) + C but the answer key says (2/9)(t^(3/2))(3ln(t) - 2) + C Sorry for lack of symbols, I know it's hard to read.

OpenStudy (anonymous):

Which did you take as your 'u' and which as your 'dv'?

OpenStudy (anonymous):

u = ln(t), du = 1/t (dt); dv = t^(1/2)dt, v = (2t^(3/2)/3)

OpenStudy (lgbasallote):

your mistake seems to be in vdu

OpenStudy (lgbasallote):

\[\huge \int vdu = \frac 23\int t^{3/2 - 2/2}\] is that what you did?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

which gets 2/3 * integral of sqrt t

OpenStudy (anonymous):

OHHHHHHHHH

OpenStudy (anonymous):

i see what i did wrong, i think i'll get the right answer, thanks

OpenStudy (lgbasallote):

welcome